# Analysis of a class of Recursive Distributional Equations Including the Resistance of the Series-parallel graph

### v1

@@ 5.2. Proof of Corollary 1
It only remains to reinterpret Theorem 2 in terms of the $\text{Beta}(2,1)$ and $\text{Beta}(2,2)$ distributions.
Let $\{X^{(n)}\}$ be the solution of the RDE (9) started from an arbitrary initial datum. According to the theorem,
(53) $\lim_{N\to\infty}\mathbb{P}\{N^{-1/\alpha}X^{(N)}\leq x\}=F(x,1),$
@@
If $\sigma<0$, then $\mathcal{F}(p,X)=\sigma|p|^{2}$ and
@@
$\lim_{N\to\infty}\mathbb{P}\{-\frac{\text{sgn}(\sigma)}{2\sqrt{|\sigma|}N^{1/2}}X^{(N)}\leq x\}=F(2\sqrt{|\sigma|}x,1)=\left\{\begin{array}[]{r l}0,&\text{if}\,\,x<0,\\ x^{2},&\text{if}\,\,0\leq x\leq 1,\\ 1,&\text{otherwise.}\end{array}\right$.
The right-hand side is precisely the CDF of the $\text{Beta}(2,1)$ distribution. Thus, case (i) is complete.
If $\sigma=0$ and $a>0$, then $\mathcal{F}(p,X)=a|p|X$ and Proposition 11 asserts that $F(x,t)=\int_{-\infty}^{x}\rho(y,t)\,dy$, where $\rho$ is the unique continuous distributional solution of the porous medium equation $\partial_{t}\rho=\frac{1}{2}a\partial_{x}^{2}(\rho^{2})$ with initial datum $\rho(0)=\delta_{0}$. This solution, called the Barenblatt (or ZKB) solution, is given by
@@
and the integrand $6y(1-y)\mathbf{1}_{[0,1]}(y)$ is nothing but the PDF of the $\text{Beta}(2,2)$ distribution. This proves case (ii).
Finally, if $\sigma=a=0$, then the formula in Section 1.3.2 implies that $(3s-g_{\rho}(s))(s-g_{\rho}(s))=0$ for all $s\geq 0$ and $\rho\in\{+,-\}$ almost surely. Since $g_{+}(s),g_{-}(s)\leq s$ hold automatically by definition, this implies that
@@
and thus $X^{(n)}\overset{d}{=}X^{(0)}$ for all $n$. This is case (iii).
@@ 5.3. Proof of Proposition 8
Let $\omega_{\Phi}$ denote the modulus of continuity of the second derivative $\partial_{x}^{2}\Phi$:
@@
$\omega_{\Phi}(\epsilon)=\sup\left\{|\partial_{x}^{2}\Phi(x)-\partial_{x}^{2}\Phi(y)|\,\mid\,x,y\in\mathbb{R},\,\,|x-y|\leq\epsilon\right\}.$
In view of the representation of $\mathscr{L}$ from Proposition 4 and symmetry (see Section 2.2), it suffices to prove that
(54) $\int_{-\infty}^{\delta^{-1}x}\mathbf{E}[\Phi_{\delta}(y_{1}+g_{+}(\delta^{-1}x-y_{1}))-\Phi_{\delta}(\delta^{-1}x)]\partial_{x}\Phi_{\delta}(y_{1})\,dy_{1}$
$\quad=-\delta^{2}(\partial_{x}\Phi(x))^{2}\int_{0}^{\infty}\mathbf{E}[|g_{+}(u)-u|]\,du$
$\qquad+\delta^{3}\Phi^{\prime}(x)\Phi^{\prime\prime}(x)\left(\frac{1}{2}\int_{0}^{\infty}\mathbf{E}[(g_{+}(u)-u)^{2}]\,du-\int_{0}^{\infty}u\mathbf{E}[g_{+}(u)-u]\,du\right)$
@@
$\qquad+C_{0}\delta^{3}\|\Phi\|_{C^{2}(\mathbb{R})}\left(\int_{0}^{\infty}\mathbf{E}[\omega_{\Phi}(\delta|g(u)-u|)\{g_{+}(u)-u\}^{2}]\,du\right$.
$\qquad\left.+\int_{0}^{\infty}u\omega_{\Phi}(\delta u)\mathbf{E}[|g(u)-u|]\,du+\delta\|\Phi\|_{C^{2}(\mathbb{R})}\int_{a}^{\infty}\mathbf{E}[\{g_{+}(u)-u\}^{2}]u\,du\right),$
where $\|\Phi\|_{C^{2}(\mathbb{R})}=\|\Phi\|_{\sup}+\|\partial_{x}\Phi\|_{\sup}+\|\partial^{2}_{x}\Phi\|_{\sup}$. To this end, first, notice that
$\int_{-\infty}^{\delta^{-1}x}\mathbf{E}[\Phi_{\delta}(y_{1}+g_{+}(\delta^{-1}x-y_{1}))-\Phi_{\delta}(\delta^{-1}x)]\partial_{x}\Phi_{\delta}(y_{1})\,dy_{1}$
$\quad=\int_{-\infty}^{x}\mathbf{E}[\Phi(x_{1}+\delta g_{+}(\delta^{-1}(x-x_{1})))-\Phi(x)]\partial_{x}\Phi(x_{1})\,dx_{1}.$
Next, it is worth noting that, by definition, the difference $s\mapsto g_{+}(s)-s$ is bounded above and below as follows
$-\sup\left\{|f(u)|\,\mid\,u\geq 0\right\}\leq g_{+}(s)-s\leq 0.$
Therefore, the assumption (15) implies that the integrals above are all finite.
Let $L$ and $Q$ be the Taylor polynomials
$L(x_{1})=\partial_{x}\Phi(x)+\partial_{x}^{2}\Phi(x)(x_{1}-x),\quad Q(x_{1})=\partial_{x}\Phi(x)(x_{1}-x)+\frac{1}{2}\partial_{x}^{2}\Phi(x)(x_{1}-x)^{2}.$
For notational ease, define $x_{1,g}$ according to the formula
$x_{1,g}=x_{1}+\delta g(\delta^{-1}[x-x_{1}]).$
The term of interest then becomes
$\int_{-\infty}^{x}\{\Phi(x_{1,g})-\Phi(x)\}\partial_{x}\Phi(x_{1})\,dx_{1}$ $=\int_{-\infty}^{x}L(x_{1})Q(x_{1,g})\,dx_{1}$
$\quad+\int_{-\infty}^{x}L(x_{1})\{\Phi(x_{1,g})-\Phi(x)-Q(x_{1,g})\}\,dx_{1}$
$\quad+\int_{-\infty}^{x}\{\partial_{x}\Phi(x_{1})-L(x_{1})\}\{\Phi(x_{1,g})-\Phi(x)\}\,dx_{1}.$
Expanding the main term and using the change-of-variables $x_{1}=x+\delta u$ leads to
$\int_{-\infty}^{x}L(x_{1})\mathbf{E}[Q(x_{1,g})]\,dx_{1}$ $=\delta^{2}(\partial_{x}\Phi(x))^{2}\int_{0}^{\infty}\mathbf{E}[g_{+}(u)-u]\,du$
@@
$\quad+\frac{1}{2}\delta^{3}\partial_{x}\Phi(x)\partial_{x}^{2}\Phi(x)\int_{-\infty}^{x}\mathbf{E}[(g_{+}(u)-u)^{2}]\,dx_{1}$
$\quad-\delta^{3}\partial_{x}\Phi(x)\partial_{x}^{2}\Phi(x)\int_{-\infty}^{x}u\mathbf{E}[g_{+}(u)-u]\,dx_{1}$
@@
$\quad-\frac{1}{2}\delta^{4}(\partial_{x}^{2}\Phi(x))^{2}\int_{-\infty}^{x}u\mathbf{E}[(g(u)-u)^{2}]\,dx_{1}.$
Notice that the order $\delta^{2}$ term has the desired form: In general, $g_{+}(s)\leq s$ always holds so $g_{+}(s)-s=-|g_{+}(s)-s|$.
It only remains to estimate the error terms:
$\left|\int_{-\infty}^{x}L(x_{1})\{\Phi(x_{1,g})-\Phi(x)-Q(x_{1,g})\}\,dx_{1}\right|$
$\quad\leq\|\Phi\|_{C^{2}(\mathbb{R})}\int_{-\infty}^{x}\omega_{\Phi}(|x_{1,g}-x|)\left(1+|x_{1}-x|\right)\{x_{1,g}-x\}^{2}\,dx_{1}$
$\quad=\delta^{3}\|\Phi\|_{C^{2}(\mathbb{R})}\int_{0}^{\infty}\omega_{\Phi}(\delta|g(u)-u|)(1+\delta|u|)\{g(u)-u\}^{2}\,du,$
@@
$\left|\int_{-\infty}^{x}\{\partial_{x}\Phi(x_{1})-L(x_{1})\}\{\Phi(x_{1,g})-\Phi(x)\}\,dx_{1}\right|$
$\quad\leq\|\Phi\|_{C^{2}(\mathbb{R})}\int_{-\infty}^{x}\omega_{\Phi}(|x_{1}-x|)|x_{1}-x||x_{1,g}-x|\,dx_{1}$
$\quad=\delta^{3}\|\Phi\|_{C^{2}(\mathbb{R})}\int_{0}^{\infty}\omega_{\Phi}(\delta u)u|g(u)-u|\,du.$

### v2

@@ 5.3. Proof of Corollary 1
It only remains to reinterpret Theorems 1 and 2 in terms of the $\text{Beta}(2,1)$ and $\text{Beta}(2,2)$ distributions.
In what follows, consider a solution $\{X^{(n)}\}$ of the RDE (9) with $\mathbf{P}\{f_{+},f_{-}\in\mathcal{S}\}=1$, (constant) bias parameter $p=1/2$, and started from an arbitrary real-valued initial datum $X^{(0)}$. Recall that assumption (15) is in effect.
Case (i): Diffusive Asymptotics. Let $\{F_{n}\}$ be the sequence of CDF’s associated with $\{X^{(n)}\}$. Define $F_{N}(x,t)=F_{[Nt]}(N^{1/2}x)$ as in Theorem 1. Since $X^{(0)}$ is real-valued,
$\lim_{N\to\infty}F_{N}(x,0)=\lim_{N\to\infty}\mathbb{P}\{N^{-1/2}X^{(0)}\leq x\}=\mathbf{1}_{[0,+\infty)}(x)\quad\text{for each}\,\,x\in\mathbb{R}\setminus\{0\}.$
First, consider the case when $\sigma\neq 0$. By Theorem 1,
(50) $\lim_{N\to\infty}\mathbb{P}\{N^{-\frac{1}{2}}X^{(N)}\leq x\}=F(x,1),$
@@
If $\sigma<0$, then $F$ is determined by the formula
@@
$\lim_{N\to\infty}\mathbb{P}\{-\frac{\text{sgn}(\sigma)}{2\sqrt{|\sigma|}N^{1/2}}X^{(N)}\leq x\}=F(2x\sqrt{|\sigma|},1)=\left\{\begin{array}[]{r l}0,&\text{if}\,\,x<0,\\ x^{2},&\text{if}\,\,0\leq x\leq 1,\\ 1,&\text{otherwise.}\end{array}\right$.
The righthand side is precisely the CDF of the $\text{Beta}(2,1)$ distribution. Therefore, $2^{-1}N^{-1/2}X^{(N)}\overset{d}{\to}\sqrt{|\sigma|}\text{Beta}(2,1)$.
It only remains to consider the case when $\sigma=0$. Fix a continuous function $G_{\text{in}}\in CDF(\overline{\mathbb{R}})$ such that $G_{\text{in}}(x)\leq\mathbf{1}_{[0+\infty)}(x)$ for each $x\in\mathbb{R}$. Consider the sequence $\{Y^{(N,n)}\}$ obtained by solving the RDE (9) with an $N$-dependent initial datum $Y^{(N,0)}$ chosen in such a way that, for every $x\in\mathbb{R}$,
(51) $\mathbb{P}\{Y^{(N,0)}\leq x\}\leq\mathbb{P}\{X^{(0)}\leq x\}\quad\text{for each}\,\,N\in\mathbb{N},$
$\lim_{N\to\infty}\mathbb{P}\{N^{-1/2}Y^{(N,0)}\leq x\}=G_{\text{in}}(x).$
This is equivalent to asking that $Y^{(N,0)}=X^{(0)}+N^{1/2}Z^{(N)}$ for some sequence of nonnegative random variables $\{Z^{(N)}\}$ with law converging to $G_{\text{in}}$.
Since $G_{\text{in}}$ is continuous, $\sigma=0$, and $p=1/2$ independently of $N$, applying Theorem 3 with $T^{(N)}=T$, $\delta^{(N)}=N^{-1/2}$, $\mathcal{F}=0$, and $Q^{(N)}=Q=0$ yields
$\lim_{N\to\infty}\mathbb{P}\{N^{-1/2}Y^{(N,N)}\leq x\}=G(x,1),$
where $G$ is the unique bounded viscosity solution of the equation $\partial_{t}G=0$ in $\mathbb{R}\times(0,+\infty)$ with initial condition $G(x,0)=G_{\text{in}}(x)$. This is just the time-independent function $G(x,t)=G_{\text{in}}(x)$. In particular, by (51) and monotonicity (Proposition 2),
$G_{\text{in}}(x)=\lim_{N\to\infty}\mathbb{P}\{N^{-1/2}Y^{(N,N)}\leq x\}\leq\liminf_{N\to\infty}\mathbb{P}\{N^{-1/2}X^{(N)}\leq x\}.$
Since $G_{\text{in}}$ was an arbitrary continuous CDF lying below $\mathbf{1}_{[0,+\infty)}$, this implies
$\mathbf{1}_{[0,+\infty)}(x)\leq\liminf_{N\to\infty}\mathbb{P}\{N^{-1/2}X^{(N)}\leq x\}\quad\text{for each}\,\,x\in\mathbb{R}\setminus\{0\}.$
Approximating the step function $\mathbf{1}_{[0,+\infty)}$ this time from above rather than below, one similarly finds
$\limsup_{N\to\infty}\mathbb{P}\{N^{-1/2}X^{(N)}\leq x\}\leq\mathbf{1}_{[0,+\infty)}(x)\quad\text{for each}\,\,x\in\mathbb{R}\setminus\{0\}.$
Therefore, $N^{-1/2}X^{(N)}\overset{d}{\to}0$ as $N\to\infty$, and case (i) is complete.
Case (ii): Subdiffusive Asymptotics. Assume the stronger condition (15) and that $\sigma=0$ and $a>0$. Since as in the previous step $N^{-1/3}X^{(0)}\to 0$ in distribution as $N\to+\infty$, Theorem 2 implies that
(52) $\lim_{N\to\infty}\mathbb{P}\{N^{-\frac{1}{3}}X^{(N)}\leq x\}=F(x,1)$
provided $F$ is the unique bounded discontinuous viscosity solution of (49) with initial datum $F_{\text{in}}=\mathbf{1}_{[0,\infty)}$. By Proposition 15 below, $F(x,t)=\int_{-\infty}^{x}\rho(y,t)\,dy$, where $\rho$ is the unique continuous distributional solution of the porous medium equation $\partial_{t}\rho=\frac{1}{2}a\partial_{x}^{2}(\rho^{2})$ with initial datum $\rho(0)=\delta_{0}$. According to [24], this solution, called the Barenblatt (or ZKB) solution, is given by
@@
and the integrand $6y(1-y)\mathbf{1}_{[0,1]}(y)$ is nothing but the PDF of the $\text{Beta}(2,2)$ distribution. This completes the proof in case (ii).
Case (iii): Triviality when $\sigma=a=0$. Finally, suppose that (16) holds and $\sigma=a=0$. Since $g_{\pm}(s)\leq s$ holds automatically by definition, the identity $a=0$ implies that $(3s-g_{\rho}(s))(s-g_{\rho}(s))=0$ for all $s\geq 0$ and $\rho\in\{+,-\}$ $\mathbf{P}$-almost surely. In particular,
@@
and thus $X^{(n)}\overset{d}{=}(1-\Theta)\max\{X^{(n-1)}_{1},X^{(n-1)}_{2}\}+\Theta\min\{X^{(n-1)}_{1},X^{(n-1)}_{2}\}\overset{d}{=}X^{(n-1)}$ for all $n$ (since $\mathbb{P}\{\Theta=1\}=p=1/2$). In other words, $\{X^{(n)}\}$ is the constant sequence $X^{(n)}\overset{d}{=}X^{(0)}$, completing the proof of case (iii). ∎
@@ 5.4. Proof of Proposition 10
It will be convenient to start by fixing some notation. Throughout this proof, define $h_{+},h_{-}:[0,+\infty)\to[0,+\infty)$ by
$h_{\pm}(s)=s-g_{\pm}(s).$
Notice that $0\leq h_{\pm}(s)\leq s$ for any $s\geq 0$ by the definition of $g_{\pm}$ (Section 1.4.1).
Define coefficients $\sigma_{+}$ and $\sigma_{-}$ by
(53) $\sigma_{+}=-\int_{0}^{+\infty}\mathbf{E}h_{+}(s)\,ds,\quad\sigma_{-}=\int_{0}^{+\infty}\mathbf{E}h_{-}(s)\,ds,$
which are finite by (15), so that $\sigma=\sigma_{+}+\sigma_{-}$. In the case of the stronger assumption (16), also define coefficients $a_{+}$ and $a_{-}$ by
$a_{\pm}=\int_{0}^{+\infty}s\mathbf{E}h_{\pm}(s)\,ds+\frac{1}{2}\int_{0}^{+\infty}\mathbf{E}h_{\pm}(s)^{2}\,ds=\frac{1}{2}\int_{0}^{+\infty}\mathbf{E}(2s+h_{\pm}(s))h_{\pm}(s)\,ds.$
By definition, $a=a_{+}+a_{-}$.
Assumption (15) says that $\int_{0}^{+\infty}\mathbf{E}h_{\pm}(s)\,ds<+\infty$, while assumption (16) says that $\int_{0}^{+\infty}(1+s)\mathbf{E}h_{\pm}(s)\,ds<+\infty$. Notice that, in the latter case, the bound $0\leq h_{\pm}(s)\leq s$ implies, in addition, that
(54) $\int_{0}^{+\infty}\mathbf{E}h_{\pm}(s)^{2}\,ds\leq\int_{0}^{+\infty}s\mathbf{E}h_{\pm}(s)\,ds<+\infty.$
In particular, $a_{+}$ and $a_{-}$ are finite.
With these preliminaries out of the way, fix a $\Phi\in CDF(\overline{\mathbb{R}})$ that has bounded, uniformly continuous first and second derivatives, and define $\Phi_{\delta}(y)=\Phi(\delta y)$ for $\delta>0$. Let $\omega_{\Phi}$ be the modulus of continuity
$\omega_{\Phi}(\delta)=\sup\left\{|\partial_{x}\Phi(x)-\partial_{x}\Phi(y)|+|\partial_{x}^{2}\Phi(x)-\partial_{x}^{2}\Phi(y)|\,\mid\,x,y\in\mathbb{R},\,\,|x-y|\leq\delta\right\}.$
Recall from Section 3 that it is possible to write
(55) $\mathscr{L}^{(N)}=2p^{(N)}\mathscr{L}_{+}+2(1-p^{(N)})\mathscr{L}_{-}$
for some $N$-independent functions $\mathscr{L}_{+}$ and $\mathscr{L}_{-}$ defined on $CDF(\overline{\mathbb{R}})$. Thus, the scaling behavior of $\mathscr{L}^{(N)}$ is determined by that of $\mathscr{L}_{+}$ and $\mathscr{L}_{-}$. Further, since the assumption (15) implies that $\mathbf{P}\{f_{+}(+\infty)=f_{-}(+\infty)=1\}$ (as proved in Section 5.2 above), the formula (33) can be invoked to find
$(\mathscr{L}_{+}\Phi_{\delta})(\delta^{-1}x)$ $=\delta\int_{0}^{+\infty}\mathbf{E}[\Phi(x-\delta h_{+}(s))-\Phi(x)]\partial_{x}\Phi(x-\delta s)\,ds,$
$(\mathscr{L}_{-}\Phi_{\delta})(\delta^{-1}x)$ $=\delta\int_{0}^{+\infty}\mathbf{E}[\Phi(x+\delta h_{-}(s))-\Phi(x)]\partial_{x}\Phi(x+\delta s)\,ds.$
In view of (55), it only remains to prove the following two claims: First, under assumption (15), for any $\rho\in\{+,-\}$,
@@
(56) $\lim_{\delta\downarrow 0}\sup\left\{\left|\delta^{-2}(\mathscr{L}_{\rho}\Phi_{\delta})(\delta^{-1}x)-\sigma_{\rho}|\partial_{x}\Phi(x)|^{2}\right|\,\mid\,x\in\mathbb{R}\right\}=0.$
Further, if assumption (16) holds, then
(57) $\lim_{\delta\downarrow 0}\sup\left\{\left|\delta^{-3}(\mathscr{L}_{\rho}\Phi_{\delta})(\delta^{-1}x)-\delta^{-1}\sigma_{\rho}|\partial_{x}\Phi(x)|^{2}-a_{\rho}|\partial_{x}\Phi(x)|\partial_{x}^{2}\Phi(x)\right|\,\mid\,x\in\mathbb{R}\right\}=0.$
$\delta^{2}$ Asymptotics. Assume (15) holds and fix a $\rho\in\{+,-\}$. First, observe that, by the definition (53) of $\sigma_{\rho}$,
(58) $\delta^{-2}(\mathscr{L}_{\rho}\Phi_{\delta})(\delta^{-1}x)-\sigma_{\rho}|\partial_{x}\Phi(x)|^{2}$
$=\mathbf{E}\int_{0}^{+\infty}\left(\frac{1}{\delta h_{\rho}(s)}\{\Phi(x-\rho\delta h_{\rho}(s))-\Phi(x)\}\partial_{x}\Phi(x-\rho\delta s)+\rho|\partial_{x}\Phi(x)|^{2}\right)h_{\rho}(s)\,ds.$
By the assumptions on $\Phi$ and the bound $h_{\rho}(s)\leq s$,
$\left|\mathbf{E}\int_{0}^{+\infty}\left(\frac{1}{\delta h_{\rho}(s)}\{\Phi(x-\rho\delta h_{\rho}(s))-\Phi(x)\}\partial_{x}\Phi(x-\rho\delta s)+\rho|\partial_{x}\Phi(x)|^{2}\right)h_{\rho}(s)\,ds\right|$
@@
$\qquad\leq\|\partial_{x}\Phi\|_{\text{sup}}\mathbf{E}\int_{0}^{+\infty}\left(\omega_{\Phi}(\delta h_{+}(s))+\omega_{\Phi}(\delta s)\right)h_{\rho}(s)\,ds$
$\qquad\leq 2\|\partial_{x}\Phi\|_{\text{sup}}\int_{0}^{+\infty}\omega_{\Phi}(\delta s)\mathbf{E}h_{\rho}(s)\,ds$
Since assumption (15) implies $\mathbf{E}h_{\rho}$ is integrable, Lebesgue’s dominated convergence theorem implies that the righthand side vanishes as $\delta\downarrow 0$. Therefore, (56) follows.
$\delta^{3}$ Asymptotics. Finally, assume that assumption (16) holds. In this case, notice that the main error term from the previous step can be rewritten in the form
$\mathbf{E}\int_{0}^{+\infty}\left(\frac{1}{\delta h_{\rho}(s)}\{\Phi(x-\rho\delta h_{\rho}(s))-\Phi(x)\}\partial_{x}\Phi(x-\rho\delta s)+\rho|\partial_{x}\Phi(x)|^{2}\right)h_{\rho}(s)\,ds$
$\quad=\rho\partial_{x}\Phi(x)\mathbf{E}\int_{0}^{+\infty}(\partial_{x}\Phi(x)-\partial_{x}\Phi(x-\rho\delta s))h_{\rho}(s)\,ds$
@@
$\quad\quad+\mathbf{E}\int_{0}^{+\infty}\left(\frac{1}{\delta h_{\rho}(s)}\{\Phi(x-\rho\delta h_{\rho}(s))-\Phi(x)\}+\rho\partial_{x}\Phi(x)\right)\partial_{x}\Phi(x-\rho\delta s)h_{\rho}(s)\,ds.$
In the first term, one finds
$\left|\frac{\rho}{\delta}\mathbf{E}\int_{0}^{+\infty}\partial_{x}\Phi(x)(\partial_{x}\Phi(x)-\partial_{x}\Phi(x-\rho\delta s))h_{\rho}(s)\,ds-\partial_{x}\Phi(x)\partial_{x}^{2}\Phi(x)\int_{0}^{+\infty}s\mathbf{E}h_{\rho}(s)\,ds\right|$
$\qquad\leq\|\partial_{x}\Phi\|_{\text{sup}}\int_{0}^{+\infty}\omega_{\Phi}(\delta s)s\mathbf{E}h_{\rho}(s)\,ds.$
Similarly, in the second term, again using $\omega_{\Phi}(\delta h_{+}(s))\leq\omega_{\Phi}(\delta s)$,
$\left|\frac{1}{\delta}\mathbf{E}\int_{0}^{+\infty}\left(\frac{1}{\delta h_{\rho}(s)}\{\Phi(x-\rho\delta h_{\rho}(s))-\Phi(x)\}+\rho\partial_{x}\Phi(x)\right)\partial_{x}\Phi(x-\rho\delta s)h_{\rho}(s)\,ds\right$.
$\qquad\left.-\frac{1}{2}\partial_{x}\Phi(x)\partial_{x}^{2}\Phi(x)\int_{0}^{+\infty}\mathbf{E}h_{\rho}(s)^{2}\,ds\right|$
@@
$\qquad\leq\frac{1}{2}\left(\|\partial_{x}\Phi\|_{\text{sup}}+\|\partial_{x}^{2}\Phi\|_{\text{sup}}\right)\int_{0}^{+\infty}\omega_{\Phi}(\delta s)\mathbf{E}h_{\rho}(s)^{2}\,ds.$
Combining these estimates with the identity (58) and estimate (54), one finds
$\sup\left\{\left|\delta^{-3}(\mathscr{L}_{\rho}\Phi_{\delta})(\delta^{-1}x)-\delta^{-1}\sigma_{\rho}|\partial_{x}\Phi(x)|^{2}-a_{\rho}|\partial_{x}\Phi(x)|\partial_{x}^{2}\Phi(x)\right|\,\mid\,x\in\mathbb{R}\right\}$
@@
$\qquad\leq\left(1+\|\partial_{x}\Phi\|_{\text{sup}}+\|\partial_{x}^{2}\Phi\|_{\text{sup}}\right)\int_{0}^{+\infty}\omega_{\Phi}(\delta s)s\mathbf{E}h_{\rho}(s)\,ds.$
Since the righthand side vanishes as $\delta\downarrow 0$ by dominated convergence, (57) follows. ∎
5.5. Counterexample
Here is a counterexample, showing that $N^{-1/2}X^{(N)}$ can blow up if (15) fails to hold. Consider the law $\mathbf{P}$ under which $(f_{+},f_{-})$ are given by
$f_{+}(u)=(Z-u)_{+},\quad f_{-}(u)=0,$
where $Z$ is a nonnegative random variable such that $\mathbb{E}Z^{2}=+\infty$. With this choice of forcing, the RDE (9) with $p=1/2$ is superdiffusive in the sense that if $\{X^{(n)}\}$ is a solution started from any initial datum and $p=1/2$, then
$N^{-\frac{1}{2}}X^{(N)}\overset{d}{\to}+\infty\quad\text{as}\,\,N\to+\infty.$
To see this, it is convenient to define $f^{(m)}_{+}$ for $m\in\mathbb{N}$ by
$f^{(m)}_{+}(u)=(\max\{Z,m\}-u)_{+}.$
Let $\{X^{(m,n)}\}$ be a solution of the RDE (9) started from the same initial distribution as $\{X^{(n)}\}$, but with the forcing $(f_{+},0)$ replaced by $(f_{+}^{(m)},0)$. Note that the inverse $g_{+}^{(m)}$ associated with $f_{+}^{(m)}$ is given by
$g_{+}^{(m)}(s)=0\quad\text{if}\,\,s<\max\{Z,m\},\quad g_{+}^{(m)}(s)=s,\quad\text{otherwise,}$
and the inverse $g_{-}$ associated with $f_{-}=0$ is $g_{-}(s)=s$. Therefore, the RDE with $(f^{(m)}_{+},0)$ satisfies (15) and the constant $\sigma^{(m)}$ determining the asymptotics is given by
$\sigma^{(m)}=-\mathbf{E}\int_{0}^{\max\{Z,m\}}s\,ds=-\frac{1}{2}\mathbf{E}\max\{Z^{2},m^{2}\}.$
Note, in particular, that $|\sigma^{(m)}|\to+\infty$ as $m\to+\infty$ by the choice of $Z$.
Since $f_{+}\geq f^{(m)}_{+}$, monotonicity with respect to the data $f_{+}$ (Proposition 3) implies that, for any $x\in\mathbb{R}$,
$\mathbb{P}\{N^{-1/2}X^{(N)}\leq x\}\leq\mathbb{P}\{N^{-1/2}X^{(m,N)}\leq x\}.$
Therefore, by Corollary 1,
$\limsup_{N\to\infty}\mathbb{P}\{N^{-1/2}X^{(m,N)}\leq x\}\leq F(2^{-1}|\sigma^{(m)}|^{-\frac{1}{2}}x).$
where $F$ is the CDF of the $\text{Beta}(2,1)$ random variable Sending $m\to+\infty$, this implies
$\limsup_{N\to\infty}\mathbb{P}\{N^{-1/2}X^{(N)}\leq x\}\leq F(0)=0\quad\text{for each}\,\,x\in\mathbb{R}.$